OpenStudy (anonymous):

Solve –4(x + 2) = 16

OpenStudy (anonymous):

-4x - 8 = 16

OpenStudy (anonymous):

I addd 8 to both sides correct?

OpenStudy (anonymous):

then divide by - 4?

ganeshie8 (ganeshie8):

thats super... go ahead. solve

ganeshie8 (ganeshie8):

how much u got finally

OpenStudy (anonymous):

- 6?

ganeshie8 (ganeshie8):

correct.

ganeshie8 (ganeshie8):

you can confirm it by putting x= -6 in original equation.. and seeing if it equals right hand side

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

Okay cool, I got it (:

OpenStudy (anonymous):

alternate: –4(x + 2) = 16 divide both side by -4 x+2=16/-4 x+2=-4 subtract both sides by 2 x=-4-2=-6 ..ans

OpenStudy (anonymous):

any question?......... hope this is much easier for u katiebae

OpenStudy (anonymous):

Thank you. Got it! :D

OpenStudy (anonymous):

ok....